# RHYTHM: Reasoning with Hierarchical Temporal Tokenization for Human Mobility

### v1

@@ Abstract
Predicting human mobility is inherently challenging due to complex long-range dependencies and multi-scale periodic behaviors. To address this, we introduce RHYTHM (Reasoning with Hierarchical Temporal Tokenization for Human Mobility), a unified framework that leverages large language models (LLMs) as general-purpose spatio-temporal predictors and trajectory reasoners. Methodologically, RHYTHM employs temporal tokenization to partition each trajectory into daily segments and encode them as discrete tokens with hierarchical attention that captures both daily and weekly dependencies, thereby significantly reducing the sequence length while preserving cyclical information. Additionally, we enrich token representations by adding pre-computed prompt embeddings for trajectory segments and prediction targets via a frozen LLM, and feeding these combined embeddings back into the LLM backbone to capture complex interdependencies. Computationally, RHYTHM freezes the pretrained LLM’s backbone to reduce attention complexity and memory cost. We evaluate our model against state-of-the-art methods using three real-world datasets. Notably, RHYTHM achieves a 2.4% improvement in overall accuracy, a 5.0% increase on weekends, and a 24.6% reduction in training time. Code is publicly available at https://github.com/he-h/rhythm.
@@ 1 Introduction
To address this challenge, we introduce RHYTHM (Reasoning with Hierarchical Temporal Tokenization for Human Mobility), a human mobility foundation model that rethinks mobility modeling via structured temporal abstraction. RHYTHM unites efficient multi-scale temporal tokenization with the deep reasoning capabilities of pretrained LLMs, delivering a scalable, accessible framework for accurate trajectory prediction with substantially reduced computational overhead.
@@
The structure of the framework is inspired by inherent patterns of human mobility. People’s movements are not random; they follow an underlying order marked by recurring patterns. Previous studies [12, 24, 18] have shown strong daily and weekly rhythms in movement patterns. Notably, Song et al. [58] quantify this regularity by showing that 93% of daily trajectories are predictable, underscoring the critical role of cyclical temporal context in mobility modeling. Capturing the complex interdependence between spatial locations and repeating time intervals—where place choices and timing influence each other in nontrivial, periodic ways—remains a central challenge. Yet existing approaches fall short: Markov and RNN-based methods either disregard long-term periodicity or suffer vanishing gradients over long sequences [76, 19, 22], while transformer-based methods treat time as static, failing to disentangle multi-scale temporal patterns [27, 72]. To bridge this gap, we decompose each trajectory into meaningful segments, tokenizing each into discrete representations that capture local patterns (e.g., morning commutes) through intra-segment attention. These segment tokens are then pooled into higher-level representations, enabling inter-segment attention to model long-range dependencies across days, as illustrated in Figure˜1, thereby reducing sequence length and quadratically lowering attention cost. To enhance semantic richness, each token is augmented with a pre-computed prompt embedding—derived from a frozen LLM encompassing trajectory context and task descriptions-before being passed into the backbone for deep reasoning.
@@
Figure 1: Motivation for RHYTHM. Instead of processing entire trajectories as a continuous sequence, RHYTHM segments trajectories into tokens to better capture periodic patterns.
@@
Recent work demonstrates LLMs’ remarkable capabilities not only as sequential representation extractors to capture the spatio-temporal attention patterns but also as reasoning models  [13, 9]. Prior works [25, 59, 17, 21] demonstrate their reasoning capabilities through techniques such as few-shot prompting [9], chain-of-thought reasoning [50, 49, 66], and in-context learning [16]. However, mobility-specific models like PMT [71] and ST-MoE-BERT [26] lack the capability to leverage LLMs for modeling complex correlations in human flows, limiting their predictive performance. By integrating an LLM-based reasoning module, RHYTHM more effectively models these complex interdependencies. To maintain scalability, RHYTHM adopts a parameter-efficient adaptation strategy by freezing the pretrained LLM and avoid extensive fine-tuning. This design captures fine-grained spatio-temporal dynamics, deep semantic context, and leverages LLM reasoning—all while minimizing computational and memory overhead—making RHYTHM ideally suited for deployment in resource-constrained, real-world environments.
@@ 3.2 Model structure
$\tilde{\mathbf{E}}^{(i)}=\mathrm{Attention}(\mathbf{s}_{i}).$
@@
$\mathbf{SE}_{i}=\mathrm{Pool}\bigl(\tilde{\mathbf{E}}^{(i)}\bigr).$
The resulting segment tokens $\{\tilde{\mathbf{SE}}_{1},\tilde{\mathbf{SE}}_{2},\ldots,\tilde{\mathbf{SE}}_{N}\}$ undergo inter-segment attention to capture broader temporal context and long-range dependencies:
$\tilde{\mathbf{SE}}_{1:N}=\operatorname{Attention}(\mathbf{SE}_{1:N}),$
yielding refined segment-level embedding $\mathbf{\tilde{SE}_{i}}\in\mathbb{R}^{D}$ that integrates contextual information across multiple temporal scales. By reducing the effective sequence length from $T$ to $N$ while preserving both fine-grained temporal dynamics and long-term dependencies, our approach addresses the computational challenges of modeling extended mobility trajectories.
@@
Cross-representational Mobility Prediction.
Since the latent space of the LLM encompasses both temporal tokens and semantic tokens, the semantic embedding can be aligned with the corresponding time span without extending the context length. Consequently, the combined embedding $\mathbf{CE}_{i}$ for segment $i$ is obtained by elementwise adding the segment embedding $\tilde{\mathbf{SE}}_{i}$ and semantic embedding $\mathbf{SE}_{i}$:
$\mathbf{CE}_{i}=\tilde{\mathbf{SE}}_{i}+\mathbf{TE}_{i}.$
Here, $\mathbf{TE}_{i}$ serves a role similar to positional embeddings [61], while avoiding the sequence length overhead incurred by prompt concatenation [32]. Similarly, the combined embedding for future timestep $T+j$ is computed as $\mathbf{CE}_{N+j}=\tilde{\mathbf{E}}_{N+j}+\mathbf{TE}_{\mathcal{T}}$, following the combined embedding for $N$ segments.
@@
$P(l_{T+j}|\mathcal{X},\mathcal{T})=\operatorname{softmax}(W_{o}\tilde{\mathbf{h}}_{N+j}+\mathbf{b}_{o}),$
where $W_{o}\in\mathbb{R}^{|\mathcal{L}|\times D}$. These logits are then used to determine the most likely location predictions, thereby generating human mobility forecasting as defined in our problem statement.
@@ 4 Experiment
To evaluate the model’s performance on mobility prediction, we use multiple pretrained LLMs as the backbone of RHYTHM. These models are obtained from Hugging Face along with their pretrained weights. The specific LLM variants used are detailed in Section˜F.4.
@@
For mobility prediction, we employ Accuracy@k, where candidate locations are ranked based on model-predicted probabilities, and a prediction is considered correct if the true location is among the top k—and Mean Reciprocal Rank (MRR) to evaluate ranking performance. These metrics have been shown to correlate well with human mobility prediction tasks [23, 19]. We also utilize Dynamic Time Warping (DTW) [47] and BLEU [51] as real-world metrics to evaluate the performance. DTW quantifies their spatial alignment, while BLEU measures the n-gram similarity between predicted and ground-truth trajectories. Detailed descriptions about the evaluation metrics can be found in Section˜F.1.
@@
We evaluate our approach on three real-world datasets collected from the cities of Kumamoto, Sapporo, and Hiroshima sourced from YJMob100K [74]. Each day is divided into 48 time slots (each representing 30 minutes), though not every slot contains an observation. Each dataset is divided into training, validation, and test sets based on days, with 70%, 20%, and 10% of the data allocated to each set, respectively. More details about the dataset can be found in Appendix˜E.
@@ Results.
Table˜1 show that RHYTHM outperforms the baselines across three datasets in most metrics. On the Sapporo and Hiroshima dataset, RHYTHM achieves the best performance in all evaluation metric. These findings underscore the effectiveness of RHYTHM in mobility prediction tasks. CMHSA and PMT may perform better in Accuracy@3 on the Kumamoto dataset due to their specialized attention mechanisms that effectively capture mid-range candidate locations in this region. Despite sharing an LLM-based architecture, Mobility-LLM underperforms compared to RHYTHM, likely because it was primarily designed for visiting intention tasks requiring rich semantic context. In contrast, RHYTHM leverages temporal tokenization and LLM to model multi-scale spatio-temporal dependencies, prioritizing precise location likelihood maximization. This design focus enables RHYTHM to excel in top-rank precision metrics. Overall, RHYTHM achieves a 2.4% improvement in Accuracy@1 and a 1.0% Accuracy@5 respectively compared to the best baseline model.
@@
We evaluate RHYTHM against baseline models using BLEU and DTW, which respectively measure n-gram similarity and spatial alignment error between predicted and ground-truth trajectories. As shown in Table˜2, RHYTHM scores the best DTW performance on Sapporo, demonstrating superior spatial alignment. While COLA leads in BLEU scores for all cities, RHYTHM ranks second in Kumamoto. This highlights a key trade-off between exact sequence matching and minimizing spatial deviations. One potential explanation is that COLA employs a post-hoc adjustment technique that recalibrates predictions to better align with the long-tail frequency distribution of locations, which may enhance mid-tier accuracy by mitigating overconfidence in dominant locations. Notably, RHYTHM significantly outperforms LSTM-based methods and transformer baselines by leveraging temporal tokenization and prompt-guided reasoning to enhance sequential coherence and spatial precision. This results in an optimal balance for real-world mobility tasks. For MRR, RHYTHM consistently outperforms all baselines, achieving a 1.44% improvement over the best baseline and demonstrates its superior ranking capability across diverse mobility patterns. Additional experimental results and extended evaluations are reported in Appendix˜G.
@@
To demonstrate that RHYTHM transfers well across pretrained LLMs, we vary the size of the pretrained backbone and train it on the mobility prediction datasets (see Table˜1 for detailed results). In our experiments, we change the size of the pretrained model in RHYTHM and test them on the mobility prediction datasets. We use the LLaMA-3.2-1B, LLaMA-3.2-3B, and Gemma-2-2B model as the pretrained backbone of RHYTHM. The results indicate that the performance of RHYTHM improves as the model size increases. Notably, the LLaMA-3.2-3B and Gemma-2-2B model outperforms the LLaMA-3.2-1B model in most metrics. This result demonstrates the performance of RHYTHM scales with LLM size and suggests that larger models may achieve even greater performance improvements on larger datasets. Note that our models are pretrained with 30 epochs. It’s plausible that the LLaMA-3.2-3B model requires more epochs to fully converge and realize its full performance potential compared with LLaMA-3.2-1B model. However, the LLaMA-3.2-3B model still achieves competitive performance compared to the LLaMA-3.2-1B model. Overall, the LLaMA-3.2-3B model demonstrated a 0.40% improvement in Acc@1 compared to the LLaMA-3.2-1B model, highlighting the scalability of RHYTHM.
@@
To evaluate the training speed of RHYTHM, we conduct experiments on the Sapporo dataset using the same training configuration. We run these experiments on a single NVIDIA A100 GPU with 40GB of memory. The results are shown in Figure˜4. RHYTHM reduces the number of trainable parameters to only 12.37% of the full model size, reflecting its parameter-efficient design. In terms of runtime, it achieves a 24.6% reduction in training time compared with the best-performing baseline, while remaining faster than most other models, being 80.6% faster than LLM-based methods on average. Although RHYTHM is slower than lightweight models such as LSTM, DeepMove, PatchTST, and iTransformer, it substantially outperforms them in accuracy. Moreover, RHYTHM maintains computational efficiency comparable to PMT, COLA and ST-MoE-BERT, despite having significantly higher parameter counts, demonstrating its parameter-efficient design and scalable architecture. Furthermore, RHYTHM’s training speed scales predictably with parameter count: LLaMA-3B is 2.2 times slower than LLaMA-1B model, while Gemma-2-2B shows a 1.9 times slowdown. A detailed breakdown of preprocessing time, storage cost, and training time across datasets is provided in Appendix˜H.
@@
Figure 5: Weekly (left) and daily (right) accuracy trends of RHYTHM and baseline models on Sapporo. These results illustrate that prediction performance fluctuates over both daily and weekly intervals.
@@ 4.2 Method Analysis
All experiments utilize LLaMA-3.2-1B as the backbone model. To evaluate our key components, we conduct ablation studies across three datasets, as shown in Table˜3. Removing temporal tokenization significantly degrades performance by 5.39%, while eliminating hierarchical attention (HA) results in a 0.90% decrease, demonstrating that structured temporal encoding is the most critical element of our framework. Regarding semantic enhancement, our findings indicate that both trajectory information and task description prompts contribute substantially to RHYTHM’s effectiveness, with their removal causing a combined performance drop of 1.82%. Task descriptions yield marginally higher impact than trajectory information, with their removal causing an additional 0.10% performance decrease compared to omitting trajectory information. Further ablation studies examining the contribution of different design choices are included in Appendix˜I.

### v2

@@ Abstract
Predicting human mobility is inherently challenging due to complex long-range dependencies and multi-scale periodic behaviors. To address this, we introduce RHYTHM (Reasoning with Hierarchical Temporal Tokenization for Human Mobility), a unified framework that leverages large language models (LLMs) as general-purpose spatio-temporal predictors and trajectory reasoners. Methodologically, RHYTHM employs temporal tokenization to partition each trajectory into daily segments and encode them as discrete tokens with hierarchical attention that captures both daily and weekly dependencies, thereby quadratically reducing the sequence length while preserving cyclical information. Additionally, we enrich token representations by adding pre-computed prompt embeddings for trajectory segments and prediction targets via a frozen LLM, and feeding these combined embeddings back into the LLM backbone to capture complex interdependencies. Computationally, RHYTHM keeps the pretrained LLM backbone frozen, yielding faster training and lower memory usage. We evaluate our model against state-of-the-art methods using three real-world datasets. Notably, RHYTHM achieves a 2.4% improvement in overall accuracy, a 5.0% increase on weekends, and a 24.6% reduction in training time. Code is publicly available at https://github.com/he-h/rhythm.
@@ 1 Introduction
To address this challenge, we introduce RHYTHM (Reasoning with Hierarchical Temporal Tokenization for Human Mobility), a human mobility foundation model that rethinks mobility modeling via structured temporal abstraction. We posit that human mobility, like language, follows compositional structures: daily routines form tokens, and weekly rhythms form higher-order syntax. RHYTHM operationalizes this analogy by tokenizing time into segments and reasoning over them with a frozen large language model (LLM). This framework unites multi-scale temporal tokenization with the reasoning capabilities of pretrained LLMs, delivering a scalable and general approach for mobility prediction with reduced computational cost.
@@
The design of RHYTHM is inspired by inherent patterns of human mobility. People’s movements are not random; they follow an underlying order marked by recurring daily and weekly rhythms [12, 24, 18]. Notably, Song et al. [58] quantify this regularity by showing that 93% of daily trajectories are predictable, underscoring the critical role of cyclical temporal context in mobility modeling. Capturing this cyclical regularity requires models that can jointly represent local behaviors (e.g., morning commutes) and global temporal dependencies (e.g., weekly routines). Yet existing approaches fall short: Markov and RNN-based methods either disregard long-term periodicity or suffer vanishing gradients over long sequences [76, 19, 22], while transformer-based methods treat time as static, failing to disentangle multi-scale temporal patterns [27, 72]. To bridge this gap, we decompose each trajectory into meaningful segments, tokenizing each into discrete representations that capture local patterns through intra-segment attention. These segment tokens are then pooled into higher-level representations, enabling inter-segment attention to model long-range dependencies across days, as illustrated in Figure˜1, thereby reducing sequence length and quadratically lowering attention cost. Each segment token is augmented with pre-computed semantic embeddings derived from a frozen LLM, enriching temporal tokens with contextual meaning before being processed by the LLM backbone for reasoning.
@@
Figure 1: Motivation for RHYTHM. Instead of processing entire trajectories as a continuous sequence, RHYTHM segments them into tokens to better capture periodic patterns.
@@
Recent advances demonstrates LLMs’ remarkable capabilities not only as sequential representation extractors to capture the spatio-temporal attention patterns but also as reasoning models  [13, 9]. Prior works [25, 59, 17, 21] demonstrate their reasoning capabilities through techniques such as few-shot prompting [9], chain-of-thought reasoning [50, 49, 66], and in-context learning [16]. However, mobility-specific models like PMT [71] and ST-MoE-BERT [26] lack the capability to leverage LLMs for modeling complex correlations in human flows, limiting their predictive performance. By integrating an LLM-based reasoning module, RHYTHM more effectively models these complex interdependencies. To maintain scalability, RHYTHM adopts a parameter-efficient adaptation strategy by freezing the pretrained LLM and avoid extensive fine-tuning. This design captures fine-grained spatio-temporal dynamics, deep semantic context, and leverages LLM reasoning—all while minimizing computational and memory overhead—making RHYTHM ideally suited for deployment in resource-constrained, real-world environments.
@@ 3.2 Model structure
$\widetilde{\mathbf{E}}^{(i)}=\mathrm{Attention}(\mathbf{s}_{i}).$
@@
$\mathbf{SE}_{i}=\mathrm{Pool}\bigl(\widetilde{\mathbf{E}}^{(i)}\bigr).$
The resulting segment tokens $\{\widetilde{\mathbf{SE}}_{1},\widetilde{\mathbf{SE}}_{2},\ldots,\widetilde{\mathbf{SE}}_{N}\}$ undergo inter-segment attention to capture broader temporal context and long-range dependencies:
$\widetilde{\mathbf{SE}}_{1:N}=\operatorname{Attention}(\mathbf{SE}_{1:N}),$
yielding refined segment-level embedding $\mathbf{\widetilde{SE}_{i}}\in\mathbb{R}^{D}$ that integrates contextual information across multiple temporal scales. By reducing the effective sequence length from $T$ to $N$ while preserving both fine-grained temporal dynamics and long-term dependencies, our approach addresses the computational challenges of modeling extended mobility trajectories.
@@
Semantic–Temporal Alignment for Mobility Prediction.
Since the latent space of the LLM encompasses both temporal tokens and semantic tokens, the semantic embedding can be aligned with the corresponding time span without extending the context length. Consequently, the combined embedding $\mathbf{CE}_{i}$ for segment $i$ is obtained by elementwise adding the segment embedding $\widetilde{\mathbf{SE}}_{i}$ and semantic embedding $\mathbf{SE}_{i}$:
$\mathbf{CE}_{i}=\widetilde{\mathbf{SE}}_{i}+\mathbf{TE}_{i}.$
Here, $\mathbf{TE}_{i}$ serves a role similar to positional embeddings [61], while avoiding the sequence length overhead incurred by prompt concatenation [32]. Similarly, the combined embedding for future timestep $T+j$ is computed as $\mathbf{CE}_{N+j}=\widetilde{\mathbf{E}}_{N+j}+\mathbf{TE}_{\mathcal{T}}$, following the combined embedding for $N$ segments.
@@
$P(l_{T+j}|\mathcal{X},\mathcal{T})=\operatorname{softmax}(W_{o}\mathbf{h}_{N+j}+\mathbf{b}_{o}),$
where $W_{o}\in\mathbb{R}^{|\mathcal{L}|\times D}$. These logits are then used to determine the most likely location predictions, thereby generating human mobility forecasting as defined in our problem statement. See Appendix˜E for implementation details.
@@ 4 Experiment
To evaluate the model’s performance on mobility prediction, we use multiple pretrained LLMs as the backbone of RHYTHM. These models are obtained from Hugging Face along with their pretrained weights. The specific LLM variants used are detailed in Section˜G.4.
@@
For mobility prediction, we employ Accuracy@k, where candidate locations are ranked based on model-predicted probabilities, and a prediction is considered correct if the true location is among the top k—and Mean Reciprocal Rank (MRR) to evaluate ranking performance. These metrics have been shown to correlate well with human mobility prediction tasks [23, 19]. We also utilize Dynamic Time Warping (DTW) [47] and BLEU [51] as real-world metrics to evaluate the performance. DTW quantifies their spatial alignment, while BLEU measures the n-gram similarity between predicted and ground-truth trajectories. Detailed descriptions about the evaluation metrics can be found in Section˜G.1.
@@
We evaluate our approach on three real-world datasets collected from the cities of Kumamoto, Sapporo, and Hiroshima sourced from YJMob100K [74]. Each day is divided into 48 time slots (each representing 30 minutes), though not every slot contains an observation. Each dataset is divided into training, validation, and test sets based on days, with 70%, 20%, and 10% of the data allocated to each set, respectively. More details about the dataset can be found in Appendix˜F.
@@ Results.
Table˜1 show that RHYTHM outperforms the baselines across three datasets in most metrics. On the Sapporo and Hiroshima dataset, RHYTHM achieves the best performance in all evaluation metric. These findings underscore the effectiveness of RHYTHM in mobility prediction tasks. CMHSA and PMT may perform better in Accuracy@3 on Kumamoto due to their specialized attention mechanisms that effectively capture mid-range candidate locations in this region. Despite sharing an LLM-based architecture, Mobility-LLM underperforms compared to RHYTHM, likely because it was primarily designed for visiting intention tasks requiring rich semantic context. In contrast, RHYTHM leverages temporal tokenization and LLM to model multi-scale spatio-temporal dependencies, prioritizing precise location likelihood maximization. This design focus enables RHYTHM to excel in top-rank precision metrics. Overall, RHYTHM achieves a 2.4% improvement in Accuracy@1 and a 1.0% Accuracy@5 respectively compared to the best baseline model.
@@
We evaluate RHYTHM against baseline models using BLEU and DTW, which respectively measure n-gram similarity and spatial alignment error between predicted and ground-truth trajectories. As shown in Table˜2, RHYTHM scores the best DTW performance on Sapporo, demonstrating superior spatial alignment. While COLA leads in BLEU scores for all cities, RHYTHM ranks second in Kumamoto. This highlights a key trade-off between exact sequence matching and minimizing spatial deviations. One potential explanation is that COLA employs a post-hoc adjustment technique that recalibrates predictions to better align with the long-tail frequency distribution of locations, which may enhance mid-tier accuracy by mitigating overconfidence in dominant locations. Notably, RHYTHM significantly outperforms LSTM-based methods and transformer baselines by leveraging temporal tokenization and prompt-guided reasoning to enhance sequential coherence and spatial precision. This results in an optimal balance for real-world mobility tasks. For MRR, RHYTHM consistently outperforms all baselines, achieving a 1.44% improvement over the best baseline and demonstrates its superior ranking capability across diverse mobility patterns. Additional experimental results and extended evaluations are reported in Appendix˜H.
@@
To demonstrate that RHYTHM transfers well across pretrained LLMs, we vary the size of the pretrained backbone and train it on the mobility prediction datasets (see Table˜1 for detailed results). In our experiments, we change the size of the pretrained model in RHYTHM and test them on the mobility prediction datasets. We use the Llama-3.2-1B, Llama-3.2-3B, and Gemma-2-2B model as the pretrained backbone of RHYTHM. The results indicate that the performance of RHYTHM improves as the model size increases. Notably, Llama-3.2-3B and Gemma-2-2B model outperforms the Llama-3.2-1B model in most metrics. This result demonstrates the performance of RHYTHM scales with LLM size and suggests that larger models may achieve even greater performance improvements on larger datasets. Note that our models are pretrained with 30 epochs. It’s plausible that Llama-3.2-3B model requires more epochs to fully converge and realize its full performance potential compared with Llama-3.2-1B model. However, Llama-3.2-3B model still achieves competitive performance compared to Llama-3.2-1B model. Overall, Llama-3.2-3B model demonstrated a 0.40% improvement in Acc@1 compared to Llama-3.2-1B model, highlighting the scalability of RHYTHM.
@@
To evaluate the training speed of RHYTHM, we conduct experiments on Sapporo using the same training configuration. We run these experiments on a single NVIDIA A100 GPU with 40GB of memory. The results are shown in Figure˜4. RHYTHM reduces the number of trainable parameters to only 12.37% of the full model size, reflecting its parameter-efficient design. In terms of runtime, it achieves a 24.6% reduction in training time compared with the best-performing baseline, while remaining faster than most other models, being 80.6% faster than LLM-based methods on average. Although RHYTHM is slower than lightweight models such as LSTM, DeepMove, PatchTST, and iTransformer, it substantially outperforms them in accuracy. Moreover, RHYTHM maintains computational efficiency comparable to PMT, COLA and ST-MoE-BERT, despite having significantly higher parameter counts, demonstrating its parameter-efficient design and scalable architecture. Furthermore, RHYTHM’s training speed scales predictably with parameter count: Llama-3B is 2.2 times slower than Llama-1B model, while Gemma-2-2B shows a 1.9 times slowdown. A detailed breakdown of preprocessing time, storage cost, and training time across datasets is provided in Appendix˜I.
@@
Figure 5: Weekly (left) and daily (right) accuracy trends of RHYTHM and baselines on Sapporo. These results illustrate that prediction performance fluctuates over both daily and weekly intervals.
@@ 4.2 Method Analysis
All experiments utilize Llama-3.2-1B as the backbone model. To evaluate our key components, we conduct ablation studies across three datasets, as shown in Table˜3. Removing temporal tokenization significantly degrades performance by 5.39%, while eliminating hierarchical attention (HA) results in a 0.90% decrease, demonstrating that structured temporal encoding is the most critical element of our framework. Regarding semantic enhancement, our findings indicate that both trajectory information and task description prompts contribute substantially to RHYTHM’s effectiveness, with their removal causing a combined performance drop of 1.82%. Task descriptions yield marginally higher impact than trajectory information, with their removal causing an additional 0.10% performance decrease compared to omitting trajectory information. Further ablation studies examining the contribution of different design choices are included in Appendix˜J.